# New Bounds on Quotient Polynomials with Applications to Exact Divisibility and Divisibility Testing of Sparse Polynomials

### v3

@@ 1.5 Proof Outline
We first note that Theorem 1.10 follows immediately from Theorem 1.8 and properties of the long division algorithm. The proof of Theorem 1.8 starts, similarly to [GR11, Theorem 2.9], by considering the discrete Fourier transform of $h$. By applying the transformation on the coefficient vector $h$ we obtain
@@
where $\omega$ is a primitive $p$th root of unity. Thus, we conclude that $\left\|{h}\right\|_{2}\leq\left|h(\theta)\right|$ for some $p$th root of unity $\theta$. Since $\left|g(\theta)\right|\cdot\left|h(\theta)\right|=\left|f(\theta)\right|\leq\left\|{f}\right\|_{1}$, we get
@@
Hence, it is enough to prove an appropriate lower bound on the value of $g$ at roots of unity in order to obtain the result. To prove such a bound we consider the restriction of $g$ to the unit circle, $\tilde{g}(x)=g(e^{ix}):[0,2\pi)\to\mathbb{C}$. We prove by induction on the sparsity of $g$ that outside the neighborhood of a small set of bad points $B(g)\subset[0,2\pi)$, $g(e^{ix})$ attains large values. Indeed, when $\deg{g}=1$ this is relatively simple to show as $g$ must be of the form $x-\alpha$ (after rescaling) and since primitive roots of unity of relatively prime orders are somewhat far from each other, we can find such prime $p$ so that the values that $g$ gets on primitive $p$th roots of unity are not too small. For higher degrees, we note that if the derivative of $g$ is not small in a large enough region then $g$ is large on most of the region. Thus, by moving further away from the bad set for the derivative of $g$, $B(g^{\prime})$ we get that on this set $g$ always attains large values. We then prove by simple pigeonhole principle there exists a prime $p$ such that all primitive $p$th roots of unity are far away from the bad set and conclude the result.
The proof of Theorem 1.12 is based on two key ideas. First, we analyze the sparsity of a quotient of a division by a linear function. To do so, we use a special case of a proposition by Lenstra [Len99], which states that when $f$ can be written as $f=f_{0}+x^{d}f_{1}$ where $d-\deg{f_{0}}\geq\Omega(\log{\left\|{f}\right\|_{0}}+\log{\left\|{f}\right\|_{\infty}})$, then any zero of $f$ in $\mathbb{Q}$ that is not $1$, $-1$ nor 00 is a zero of both $f_{0}$ and $f_{1}$. We then observe that any binomial can be written as $g=x^{k}\ell(x^{m})$ for some linear function $\ell$ and $k,m\in\mathbb{N}$, and exploit this structure to conclude the bound on the sparsity of the quotient.
The proof of Theorem 1.13 is based on the same proposition by Lenstra for general roots in the algebraic closure of $\mathbb{Q}$. First, we prove a bound of the form $O\left(\left\|{f}\right\|_{0}\cdot\left(\frac{\deg{f}}{\deg{g}}\right)^{2}\right)$ on the sparsity of $h$ when $g$ is a trinomial, using an approach similar to [GGPdC21, Lemma 3.4]. We observe that trinomials have a large gap, either between the highest and second highest exponents or between the lowest and second lowest exponents. Thus, by analyzing $f/g$ over the ring of power series $\mathbb{F}[[x]]$, we conclude the result. Finally, we extend Lenstra’s approach to show that if $g$ is cyclotomic-free and $f$ can be split into two polynomials with a large degree gap, then $g$ divides $f$ if and only if $g$ divides both polynomials. This allows us to conclude that $h$ can be written as the sum of multiple quotients of polynomials whose degree ratio is smaller than $\deg{f}/\deg{g}$.
We leverage our bounds to conclude Theorem 1.14 and Theorem 1.15 by running the long division algorithm for polynomials with early termination in the case that the sparsity obtained exceeds the bound.
We prove Theorem 1.16 over finite fields by observing that for a binomial $g$, the remainder $f\bmod g$ must be sparse, and can be calculated explicitly. We prove Theorem 1.18 by first using a gap lemma for the case where $\deg{f}\leq\deg{g}(1+\frac{1}{\left\|{g}\right\|_{0}})$. In that case, $g$ must have a gap that is larger than $\deg{f}-\deg{g}$. This allows us to split $g$ into two parts, such that one of them has to be a binomial. We show that testing divisibility by this binomial is enough. Finally, we prove a reduction from testing the divisibility of a high-degree polynomial over a finite field to testing the divisibility of a polynomial of a higher sparsity but of a smaller degree, which yields a good trade-off when the field size is small. This allows us to finish the proof by showing that the case when $\deg{f}=\tilde{O}(\deg{g})$ can be reduced to the case $\deg{f}\leq\deg{g}(1+\frac{1}{\left\|{g}\right\|_{0}})$.
1.6 Open Problems
The main challenge is to extend our result over $\mathbb{Z}$ beyond trinomials and provide a polynomial time algorithm for every cyclotomic-free polynomial of bounded sparsity.
Another important question is to improve our result over $\mathbb{Z}$ to handle trinomials with cyclotomic factors, or to prove that this case is hard. We remark that our approach fails to do so only when $\deg{f}\geq\Omega(\deg{g}\log^{\omega(1)}\deg{g})$.
Another problem is extending our results over $\mathbb{F}_{p}$. For example, solving the case of $\deg{f}\geq\Omega(\deg^{1+\epsilon}{g})$ for pentanomials, extending the result to all bounded sparsities when $\deg{f}=\tilde{O}(\deg{g})$, or proving that either of those is hard are possible research directions.
Finally, we also note that the general question of the hardness of divisibility testing of lacunary polynomials (of variable sparsity) remains open, over both $\mathbb{Z}$ and $\mathbb{F}_{p}$.
1.7 Organization of the Paper
In Section 2 we set our notations and state useful facts and claims. In Section 3 we prove Theorem 1.8. In Section 4 we prove Theorems 1.12 and 1.13. In Section 5, we use the above bounds to prove Theorems 1.14 and 1.15. Finally, in Section 6 we deal with divisibility testing over finite fields and prove Theorems 1.16 and 1.18.
@@ 2.1 General Notation
Let $f=\sum_{i}f_{i}x^{i}$,. We shall use the following notation: $\left\|{f}\right\|_{0}$ denotes the sparsity of $f$ (the number of non-zero $f_{i}$’s); $\left\|{f}\right\|_{p}=\left(\sum_{i}{\left|f_{i}\right|^{p}}\right)^{1/p}$; if $R=\mathbb{Z}$ then $\left\|{f}\right\|_{\infty}=\max_{i}{\left|f_{i}\right|}$ is the height of $f$; $\text{ord}_{\alpha}{f}$ denotes the multiplicity of a root $\alpha$ in $f$, in particular, $\text{ord}_{0}{f}=\max{\left\{i:x^{i}\mid f\right\}}$; $\text{LC}(f)\in R$ denotes the leading coefficient in $f$, i.e., the coefficient of the highest degree term in $f$.
For any real or complex-valued function $f$, we define $V(f)$ to be the set of zeros of $f$, i.e., $V(f)=\left\{\alpha\mid f(\alpha)=0\right\}$. For a complex-valued function $f$ we denote by $\Re{f}$ and $\Im{f}$ the real and imaginary parts of $f$, respectively. We denote the derivative of $f$ by $f^{\prime}$.
A polynomial with $t$ non-zero terms is called a $t$-nomial; we call a polynomial with two terms a binomial, a polynomial with three terms a trinomial, and a polynomial with five terms a pentanomial.
For an integer $k\in\mathbb{N}$, we denote $[k]=\left\{1,\dots,k\right\}$. Throughout this paper, $\log$ refers to the natural logarithm. We use $\pi$ to denote the mathematical constant, while $\pi(\cdot)$ refers to the prime counting function. We use $i\;\equiv_{m}\;j$ to denote that $i=j\mod m$.
@@ 2.2 Useful Facts
2.2.1 Polynomials
An important class of polynomials, which contains the hardest instances for many problems concerning sparse polynomials, including all examples of NP-hard problems, is the class of cyclotomic polynomials. An $n$th root of unity, for a positive integer $n$, is any $\omega\in\mathbb{C}$ satisfying the equation $\omega^{n}=1$. It is said to be primitive if it is not a root of unity of any smaller order $m<n$. The minimal polynomials over $\mathbb{Q}$ of roots of unity are called cyclotomic polynomials. More explicitly, the $n$-th cyclotomic polynomial is defined as
$\Phi_{n}=\prod_{\begin{subarray}{c}1\leq k\leq n\\ \gcd{(k,n)}=1\end{subarray}}{(x-e^{2\pi i\frac{k}{n}})}=\prod_{\zeta}{(x-\zeta)},$
where $\zeta$ ranges over the primitive $n$th roots of unity. It is not hard to show that cyclotomic polynomials have integer coefficients. A cyclotomic-free polynomial is a polynomial that has no cyclotomic factors, and thus none of its zeros is a root of unity.
We next list some basic facts regarding polynomials.
Fact 2.1.
Let $R$ be a UFD, and let $f,g\in R[x]$. Then,
  1. 1.
    $g(x)\mid f(x)\iff g(x^{m})\mid f(x^{m})$ for any $0\neq m\in\mathbb{N}$.
  2. 2.
    $g\mid f\iff gh\mid fh$ for any $0\neq h\in R[x]$.
  3. 3.
    $g\mid f\iff g\mid fh$ for any $h\in R[x]$ such that $\gcd{\left\{g,h\right\}}=1$.
  4. 4.
    If $R$ is a field, then $g\mid f\iff g\mid af\iff ag\mid f$ for any $0\neq a\in R$.
Fact 2.2 (Polynomial Remainder Theorem).
Let $R$ be a commutative ring with a unit. Let $f\in R[x]$ and $a\in R$. Then $(x-a)\mid f\iff f(a)=0$.
Claim 2.3.
Let $R$ be a UFD. Let $g\in R[x]$ be a polynomial of sparsity $s$ and let $n\in\mathbb{N}$. Then $\left\|{g^{n}}\right\|_{0}\leq\binom{n+s-1}{s-1}$.
Proof.
Let $g=\sum_{i=1}^{s}g_{i}x^{d_{i}}$. Each monomial in $g^{n}$ is of the form $x^{\sum_{i=1}^{s}a_{i}d_{i}}$ where $\sum_{i=1}^{s}a_{i}=n$ and $a_{i}\geq 0$. Hence, each monomial in $g^{n}$ corresponds to at least one solution to the equation
$\sum_{i\in[s]}{a_{i}}=n$
over the non-negative integer variables $a_{1},\dots,a_{s}$, and there are exactly $\binom{n+s-1}{s-1}$ such solutions. ∎
Fact 2.4.
Let $g\in\mathbb{F}_{p}[x]$. Then $g^{p}=g(x^{p})$.
@@ 2.2.2 Complex Analysis
Since $f$ is analytic, $\Re{f}$ is harmonic. 2.5 implies that $\Re{f}$ does not vanish on the unit circle. Thus, from the irreducibility of $x^{2}+y^{2}-1$, we deduce that $V(u)$ and $V(x^{2}+y^{2}-1)$ have no component in common. From Bézout’s theorem (see [CLO13, Chapter 8.7,Theorem 10]) we conclude that
@@
where the last inequality is due to the fact that $1-\cos$ is positive and concave on both $[\frac{\pi}{2},\frac{3\pi}{2}]$ and $[-\frac{3\pi}{2},-\frac{\pi}{2}]$. ∎
@@ 2.2.3 Number Theory
Finally, we shall also use the following upper bound on binomial coefficients.
Claim 2.10.
Let $m,n\in\mathbb{N}$. Then, $\binom{n+m}{m}\leq(n+1)^{m}$.
Proof.
We prove this bound by induction on $n+m$. For the base case, assume $n+m=1$. Then, $\binom{n+m}{m}=1$ and the bound is trivial. Now assume that the bound is true for $n+m-1$. Then by Pascal’s triangle,
$\binom{n+m}{m}$ $=\binom{n+m-1}{m}+\binom{n+m-1}{m-1}$
$\leq n^{m}+(n+1)^{m-1}$
$\leq(n+1)^{m-1}n+(n+1)^{m-1}$
$=(n+1)^{m},$
which proves the inductive step. ∎
2.3 Preprocessing
In what follows, we assume that
  1. 1.
    $\text{ord}_{0}{g}=0$,
  2. 2.
    $\left\|{g}\right\|_{0}\geq 2$,
  3. 3.
    $\deg{g}\leq\deg{f}$, and
  4. 4.
    if $R$ is a field, then we also assume that $g$ and $f$ are monic.
To satisfy these assumptions, we need to introduce a preprocessing procedure to our algorithms. In particular, we can assume $\text{ord}_{0}{g}\leq\text{ord}_{0}{f}$ since $\text{ord}_{0}{g}>\text{ord}_{0}{f}$ implies that $g\nmid f$. Thus, we can divide both $f$ and $g$ by $x^{\text{ord}_{0}{g}}$ (see 2.1) and assume that $\text{ord}_{0}{g}=0$. If $g$ is a monomial, then $g\in R$, and so we can easily test $g\mid f$. If $\deg{g}>\deg{f}$ then $g\nmid f$. Finally, if $R$ is a field, the divisibility of $f$ by $g$ is invariant to multiplication by field elements (see 2.1). Thus we may assume that $g$ and $f$ are monic.
3 A Bound on the $\ell_{2}$-norm of the Quotient Polynomial
@@
where $p$ is a prime such that $p>2\deg{f}$.
@@
where $\omega\in\mathbb{C}$ is a primitive $p$th root of unity. By applying the $p\times p$ DFT to the coefficient vector of $h$, we conclude that
$\left\|{h}\right\|_{2}^{2}=\sum_{0\leq i\leq\deg{h}}{\left|h_{i}\right|^{2}}=\frac{1}{p}\sum_{0\leq i<p}{\left|h(\omega^{i})\right|^{2}}$
where the second equality follows from the fact that the DFT matrix is unitary, and so it preserves $\ell_{2}$ norms. Observe that
@@
where the second inequality is due to Cauchy-Schwartz inequality. It follows that
$\frac{1}{p}\sum_{0<i<p}{\left|h(\omega^{i})\right|^{2}}=\left\|{h}\right\|_{2}^{2}-\frac{\left|h(\omega^{0})\right|^{2}}{p}\geq\left\|{h}\right\|_{2}^{2}-\frac{\deg{f}}{p}\left\|{h}\right\|_{2}^{2}>\frac{1}{2}\left\|{h}\right\|_{2}^{2},$
where the last inequality follows from the choice of $p$. In other words, the average value of $\left|h(\omega^{i})\right|^{2}$ for $0<i<p$ is larger than $\frac{1}{2}\left\|{h}\right\|_{2}^{2}$. Hence, there exists a $p$th root of unity $\theta\neq 1$ such that
@@
Since $f=gh$ we conclude that $\left|f(\theta)\right|=\left|g(\theta)\right|\left|h(\theta)\right|$. If $g(\theta)=0$, then the claim trivially holds. As $\left|\theta\right|=1$, $\left|f(\theta)\right|\leq\left\|{f}\right\|_{1}$. Hence,
@@
To prove Theorem 1.8 we first study the case where $g$ is linear. We show that, in this case, there exists a point $\tilde{\alpha}$ on the unit circle such that $g$ attains large values on any point on the unit circle that is far enough from $\tilde{\alpha}$. To conclude, we prove that there exists a prime $p$ such that every $p$th root of unity is far from $\tilde{\alpha}$.
@@
Let $\alpha_{1},\dots,\alpha_{k}\in\mathbb{R}$. Then, any set of $k+1$ prime numbers $P$ contains $p\in P$ such that $\left|a/p-\alpha_{i}\right|\geq\frac{1}{2\left(\max{P}\right)^{2}}$ for every integer $0<\left|a\right|<p$ and $i\in[k]$.
@@
Assume for the sake of contradiction that the statement is false. Thus, for every $p\in P$ there exists $i\in[k]$ such that $\left|a/p-\alpha_{i}\right|<\frac{1}{2\left(\max{P}\right)^{2}}$ for some $0<a<p$. By the Pigeonhole Principle, there exists $i\in[k]$ such that for two different primes $p,q\in P$, $0<a<p$ and $0<b<q$ we have
$\left|a/p-\alpha_{i}\right|<\frac{1}{2\left(\max{P}\right)^{2}}$    and $\left|b/q-\alpha_{i}\right|<\frac{1}{2\left(\max{P}\right)^{2}}.$
@@
$\frac{1}{\left(\max{P}\right)^{2}}>\left|\frac{a}{p}-\alpha_{i}\right|+\left|\alpha_{i}-\frac{b}{q}\right|\geq\left|\frac{a}{p}-\frac{b}{q}\right|=\left|\frac{ap-bq}{pq}\right|\geq 1/pq>\frac{1}{\left(\max{P}\right)^{2}}.$
@@
for any prime $p$ such that $p>2\deg{f}$. We prove that there exists a suitable $p$ such that $\left|\theta-\alpha\right|$ in Equation 3.4 must be at least $(64\deg^{2}{f})^{-1}$. First, observe that if $\alpha$ is $(64\deg^{2}{f})^{-1}$-far from any point on the unit circle, then we are done. So, assume that there exists a point $\tilde{\alpha}$ on the unit circle such that $\left|\alpha-\tilde{\alpha}\right|<(64\deg^{2}{f})^{-1}$, and define $c\in[-\pi,\pi)$ such that $e^{ic}=\tilde{\alpha}$.
2.9 implies that the interval $(2\deg{f},8\deg{f}]$ contains at least 2 primes. Thus, by 3.2, there exists a prime $p$ in this interval such that
$\left|\frac{2\pi a}{p}-c\right|=2\pi\left|\frac{a}{p}-\frac{c}{2\pi}\right|\geq\frac{\pi}{64\deg^{2}{f}},$
for any integer $0<\left|a\right|<p$. Observe that by the periodicity of $e^{ix}$, any $p$th root of unity $\theta\neq 1$ can be expressed as $\theta=e^{i\frac{2\pi a}{p}}$ such that $a\in\mathbb{Z}$, $0<\left|a\right|<p$ and $\left|\frac{2\pi a}{p}-c\right|\leq\pi$. Hence, by 2.7 it holds that
$\left|\theta-\tilde{\alpha}\right|\geq\frac{2}{\pi}\left|\frac{2\pi a}{p}-c\right|\geq\frac{1}{32\deg^{2}{f}},$
@@
$\left|\theta-\alpha\right|\geq\left|\theta-\tilde{\alpha}\right|-\left|\tilde{\alpha}-\alpha\right|\geq\frac{1}{64\deg^{2}{f}}.$
@@
To generalize the above proposition, we need to establish a lower bound on $\left|g(\omega)\right|$ where $\omega\neq 1$ is a $p$th root of unity. We next show that $g$ attains large values everywhere, with the possible exception of a small neighborhood of a small number of points.
@@
Let $f:\mathbb{R}\to\mathbb{R}$ be a continuously differentiable function, $\delta>0$ and an interval $I=[a,b]$ in which each of $f$ and $f^{\prime}$ is either non-negative or non-positive. Then $|f(\beta)|\geq\delta\min_{\alpha\in I}|f^{\prime}(\alpha)|$ for every $\beta\in[a+\delta,b-\delta]$.
@@
[…]
$\geq(\beta-a)\min_{\alpha\in I}|f^{\prime}(\alpha)|$
$\geq\delta\min_{\alpha\in I}|f^{\prime}(\alpha)|.$
@@
$f(\beta)\geq f(\beta)-f(b)$ $=\int_{\beta}^{b}{-f^{\prime}(y)dy}$
$\geq(b-\beta)\min_{\alpha\in I}|f^{\prime}(\alpha)|$
$\geq\delta\min_{\alpha\in I}|f^{\prime}(\alpha)|,$

### v4

@@ 1.3 Proof Outline
We first note that Theorem 1.13 follows immediately from Theorem 1.8 and properties of the polynomial division algorithm. The proof of Theorem 1.8 starts, similarly to [GR11, Theorem 2.9], by considering the discrete Fourier transform of the coefficient vector of $h$. Parseval’s identity gives
@@
where $\omega$ is a primitive $p$th root of unity for a prime $p>\deg(h)$. Thus, we conclude that $\left\|{h}\right\|_{2}\leq\left|h(\theta)\right|$ for some $p$th root of unity $\theta$. Since $\left|g(\theta)\right|\cdot\left|h(\theta)\right|=\left|f(\theta)\right|\leq\left\|{f}\right\|_{1}$, we conclude that
@@
Hence, it is enough to lower bound the values of $g$ at roots of unity in order to upper bound $\|h\|_{2}$. To prove such a bound, we consider the restriction of $g$ to the unit circle, $\tilde{g}(x)=g(e^{ix}):[0,2\pi)\to\mathbb{C}$. We prove by induction on the sparsity of $g$ that outside the neighborhood of a small set of bad points $B(g)\subset[0,2\pi)$, $g(e^{ix})$ attains large values. Indeed, when $\deg{g}=1$ this is relatively simple to show as $g$ must be of the form $x-\alpha$ (after rescaling) and since primitive roots of unity of relatively prime orders are somewhat far from each other, we can find such prime $p$ so that the values of $g$ on primitive $p$th roots of unity are not too small. For higher degrees, we note that if the derivative of $g$ is not small in a large enough region then $g$ is large on most of that region. Thus, by moving further away from the bad set for the derivative of $g$, $B(g^{\prime})$ we get that on this set $g$ always attains large values. We then prove by simple pigeonhole principle there exists a prime $p$ such that all primitive $p$th roots of unity are far away from the bad set and conclude the result.
@@ 2.1 General Notation
Let $f=\sum_{i}f_{i}x^{i}$,. We shall use the following notation: $\left\|{f}\right\|_{0}$ denotes the sparsity of $f$ (the number of non-zero $f_{i}$’s); $\left\|{f}\right\|_{p}=\left(\sum_{i}{\left|f_{i}\right|^{p}}\right)^{1/p}$; if $R\subseteq\mathbb{C}$ then $\left\|{f}\right\|_{\infty}=\max_{i}{\left|f_{i}\right|}$ is the height of $f$; $\text{ord}_{\alpha}{f}$ denotes the multiplicity of a root $\alpha$ in $f$, in particular, $\text{ord}_{0}{f}=\max{\left\{i:x^{i}\mid f\right\}}$; $\text{LC}(f)\in R$ denotes the leading coefficient in $f$, i.e., the coefficient of the highest degree term in $f$. Similarly, $\text{TC}(f)\in R$, the trailing coefficient of $f$, denotes the coefficient of the smallest degree term in $f$.
For a polynomial $g$ we denote $g_{0}:=(g/x^{\text{ord}_{0}{g}})$ and $g_{\text{rev}}:=x^{\deg{g}}\cdot g(1/x)$. For example, if $g=\sum_{i=1}^{s}c_{i}\cdot x^{n_{i}}$ then $g_{0}=\sum_{i=1}^{s}c_{i}\cdot x^{n_{i}-n_{1}}$ and $g_{\text{rev}}=\sum_{i=1}^{s}c_{i}\cdot x^{n_{s}-n_{i}}$. Observe that $(g_{\text{rev}})_{0}=(g_{0})_{\text{rev}}$ and that $g_{0}=(g_{\text{rev}})_{\text{rev}}$.
We denote the derivative of $g$ by $g^{\prime}$.
For any real or complex-valued function $f$, we define $V(f)$ to be the set of zeros of $f$, i.e., $V(f)=\left\{\alpha\mid f(\alpha)=0\right\}$, over the corresponding field. For a complex-valued function $f$ we denote by $\Re{f}$ and $\Im{f}$ the real and imaginary parts of $f$, respectively.
For an integer $k\in\mathbb{N}$, we denote $[k]=\left\{1,\dots,k\right\}$. Throughout this paper, $\log$ refers to the natural logarithm. We use $\pi$ to denote the mathematical constant, while $\pi(\cdot)$ refers to the prime counting function.
@@ 2.2.1 Complex Analysis
Since $f$ is analytic, $\Re{f}$ is harmonic. 2.1 implies that $\Re{f}$ does not vanish on the unit circle. Thus, from the irreducibility of $x^{2}+y^{2}-1$, we deduce that $V(u)$ and $V(x^{2}+y^{2}-1)$ have no component in common. From Bézout’s theorem (see e.g., [Har92, Theorem 18.3]) we conclude that
@@
where, as before, the last inequality is due to the fact that $1-\cos$ is positive and concave on both $[\frac{\pi}{2},\frac{3\pi}{2}]$ and $[-\frac{3\pi}{2},-\frac{\pi}{2}]$. ∎
@@ 2.2.2 Number Theory
3 A Bound on the $\ell_{2}$-norm of the Factor Polynomial
@@
where $p$ is a prime such that $p>2\deg{f}$ and $g$ does not vanish at any primitive root of unity of order $p$.
@@
where $\omega\in\mathbb{C}$ is a primitive $p$th root of unity. Since the DFT matrix is unitary, applying it to the coefficient vector of $h$ we conclude that
$\left\|{h}\right\|_{2}^{2}=\sum_{0\leq i\leq\deg{h}}{\left|h_{i}\right|^{2}}=\frac{1}{p}\sum_{0\leq i<p}{\left|h(\omega^{i})\right|^{2}}.$
From
@@
we obtain, by the choice of $p$,
$\frac{1}{p}\sum_{0<i<p}{\left|h(\omega^{i})\right|^{2}}=\left\|{h}\right\|_{2}^{2}-\frac{\left|h(\omega^{0})\right|^{2}}{p}\geq\left\|{h}\right\|_{2}^{2}-\frac{\deg{f}}{p}\left\|{h}\right\|_{2}^{2}>\frac{1}{2}\left\|{h}\right\|_{2}^{2}.$
In other words, the average value of $\left|h(\omega^{i})\right|^{2}$ for $0<i<p$ is larger than $\frac{1}{2}\left\|{h}\right\|_{2}^{2}$. Hence, there exists a $p$th root of unity $\theta\neq 1$ such that
@@
Since $f=gh$ we conclude that $\left|f(\theta)\right|=\left|g(\theta)\right|\left|h(\theta)\right|$. As $\left|\theta\right|=1$, $\left|f(\theta)\right|\leq\left\|{f}\right\|_{1}$. Hence,
@@
We first consider the special case where $g$ is linear. We show that there exists a point $\tilde{\alpha}$ on the unit circle such that $g$ attains large values on any point on the unit circle that is far enough from $\tilde{\alpha}$. To conclude, we prove that there exists a prime $p$ such that every $p$th root of unity is far from $\tilde{\alpha}$.
@@
Let $\alpha_{1},\dots,\alpha_{k}\in\mathbb{R}$. Then, any set of $k+1$ prime numbers $P$ contains $p\in P$ such that $\left|a/p-\alpha_{i}\right|>\frac{1}{2\left(\max{P}\right)^{2}}$ for every integer $0<\left|a\right|<p$ and $i\in[k]$.
@@
Assume for the sake of contradiction that the statement is false. Thus, for every $p\in P$ there exists $i\in[k]$ such that $\left|a/p-\alpha_{i}\right|\leq\frac{1}{2\left(\max{P}\right)^{2}}$ for some $0<a<p$. By the Pigeonhole Principle, there exists $i\in[k]$ such that for two different primes $p,q\in P$, $0<a<p$ and $0<b<q$ we have
$\left|a/p-\alpha_{i}\right|\leq\frac{1}{2\left(\max{P}\right)^{2}}$    and $\left|b/q-\alpha_{i}\right|\leq\frac{1}{2\left(\max{P}\right)^{2}}.$
@@
$\frac{1}{\left(\max{P}\right)^{2}}\geq\left|\frac{a}{p}-\alpha_{i}\right|+\left|\alpha_{i}-\frac{b}{q}\right|\geq\left|\frac{a}{p}-\frac{b}{q}\right|=\left|\frac{ap-bq}{pq}\right|\geq 1/pq>\frac{1}{\left(\max{P}\right)^{2}}.\qed$
@@
for any prime $p$ such that $p>2\deg{f}$ and $\alpha$ is not a primitive root of unity of order $p$. We prove that there exists a suitable $p$ such that $\left|\theta-\alpha\right|$ in Equation 3.4 must be at least $(64\deg^{2}{f})^{-1}$. First, observe that if $\alpha$ is $(64\deg^{2}{f})^{-1}$-far from any point on the unit circle, then we are done. So, assume that there exists a point $\tilde{\alpha}$ on the unit circle such that $\left|\alpha-\tilde{\alpha}\right|<(64\deg^{2}{f})^{-1}$, and let $c\in[-\pi,\pi)$ such that $e^{ic}=\tilde{\alpha}$.
By 2.5, the interval $(2\deg{f},8\deg{f}]$ contains at least 2 primes. Thus, by 3.2, there exists a prime $p$ in this interval such that
$\left|\frac{2\pi a}{p}-c\right|=2\pi\left|\frac{a}{p}-\frac{c}{2\pi}\right|>\frac{\pi}{64\deg^{2}{f}},$
for any integer $0<\left|a\right|<p$. By the periodicity of $e^{ix}$, any $p$th root of unity $\theta\neq 1$ can be expressed as $\theta=e^{i\frac{2\pi a}{p}}$ such that $a\in\mathbb{Z}$, $0<\left|a\right|<p$ and $\left|\frac{2\pi a}{p}-c\right|\leq\pi$. Hence, by 2.3 it holds that
$\left|\theta-\tilde{\alpha}\right|\geq\frac{2}{\pi}\left|\frac{2\pi a}{p}-c\right|>\frac{1}{32\deg^{2}{f}},$
@@
$\left|\theta-\alpha\right|\geq\left|\theta-\tilde{\alpha}\right|-\left|\tilde{\alpha}-\alpha\right|>\frac{1}{64\deg^{2}{f}}.$
@@
We prove Theorem 1.8 in its generality by induction on $\|g\|_{0}$. Similarly to the proof outline of 3.3, we will show that $g$ attains large values everywhere, with the possible exception of a small neighborhood of a small number of points. We start with two simple claims.
@@
Let $f:\mathbb{R}\to\mathbb{R}$ be a continuously differentiable function, and $\delta>0$. Let $I=[a,b]$ be an interval in which both $f$ and $f^{\prime}$ do not change sign. That is, each of them is either non-negative or non-positive in $I$. Then, $|f(\beta)|\geq\delta\min_{\alpha\in I}|f^{\prime}(\alpha)|$ for every $\beta\in[a+\delta,b-\delta]$.
@@
$f(\beta)\geq f(\beta)-f(a)=\int_{a}^{\beta}{f^{\prime}(y)dy}\geq(\beta-a)\min_{\alpha\in I}f^{\prime}(\alpha\geq\delta\min_{\alpha\in I}|f^{\prime}(\alpha)|.$
@@
$f(\beta)\geq f(\beta)-f(b)=\int_{\beta}^{b}{-f^{\prime}(y)dy}\geq(b-\beta)\min_{\alpha\in I}|f^{\prime}(\alpha)|\geq\delta\min_{\alpha\in I}|f^{\prime}(\alpha)|,$